# PoliCon: Evaluating LLMs on Achieving Diverse Political Consensus Objectives

### v2

@@ Abstract
Achieving political consensus is crucial yet challenging for the effective functioning of social governance. However, although frontier AI systems represented by large language models (LLMs) have developed rapidly in recent years, their capabilities in this scope are still understudied. In this paper, we introduce PoliCon, a novel benchmark constructed from 2,225 high-quality deliberation records of the European Parliament over 13 years, ranging from 2009 to 2022, to evaluate the ability of LLMs to draft consensus resolutions based on divergent party positions under varying collective decision-making contexts and political requirements. Specifically, PoliCon incorporates four factors to build each task environment for finding different political consensus: specific political issues, political goals, participating parties, and power structures based on seat distribution. We also developed an evaluation framework based on social choice theory for PoliCon, which simulates the real voting outcomes of different political parties to assess whether LLM-generated resolutions meet the requirements of the predetermined political consensus. Our experimental results demonstrate that even state-of-the-art models remain undersatisfied with complex tasks like passing resolutions by a two-thirds majority and addressing security issues, while uncovering their inherent partisan biases and revealing some behaviors LLMs show to achieve the consensus, such as prioritizing the stance of the dominant party instead of uniting smaller parties, which highlights PoliCon’s promise as an effective platform for studying LLMs’ ability to promote political consensus.
@@ 1 Introduction
One of the fundamental prerequisites for effective social governance is establishing political consensus across diverse stakeholders [Prothro and Grigg, 1960, Lijphart et al., 1999, Huckfeldt et al., 2004, Rawls, 2020]. From infrastructure development to welfare policies, consensus-building underpins the legitimacy [Cohen, 2005] and implementation of collective decisions [Citrin, 2001, Potapchuk and Crocker, 2017, Shehu, 2017]. Yet, in pluralistic societies, conflicting values, power dynamics, and issue complexity render this process exceptionally challenging [Raiffa, 1982, Ehtamo et al., 1999, Susskind et al., 1999, Baker and Azher, 2024]. While large language models (LLMs) have shown promise in facilitating group discussions [Chiang et al., 2024], supporting democratic deliberation [Small et al., 2023, Fish et al., 2023, Tessler et al., 2024, Jarrett et al., 2025], resolving regional conflicts [Konya et al., 2025], and analyzing ideological stances [Chen et al., 2024, Kim et al., 2025], their capacity to find consensus in real and complex political scenarios remains underexplored. This gap raises a critical question: Can LLMs bridge divides among divergent stakeholders and achieve the objectives of different political consensus in real-world settings?
@@
Specifically, PoliCon has designed four adjustable factors to build different task environments, which are: (1) Political issues: the political problems to be discussed and their topic classification, (2) Political goals: the criteria for achieving the corresponding political consensus, (3) Participating parties: different numbers of involving stakeholders with varying stances, and (4) Power structures: differences in influence and discourse power of each party due to their number of seats. By combining these settings, we have constructed a total of 28,620 detailed scenarios. To assess whether LLM-generated resolutions meet the corresponding political goals, we further developed an open-ended evaluation framework in PoliCon based on the social choice theory [Sen, 1986, Kelly, 2013]. Through our experiments, we have verified its strong capability to simulate the real voting results, thereby allowing the effective evaluation in PoliCon (subsection 4.1).
@@ 2 Related Work
Achieving political consensus, due to its realistic and complex scenarios, the conflict of stances and values, and the need to consider diverse power structures, differs from existing works that primarily consider conversational grounding [Udagawa and Aizawa, 2019, 2020, 2021, Mitsuda et al., 2022, Mohapatra et al., 2024] and game-theoretic bargaining [Lewis et al., 2017, Paquette et al., 2019, Zhou et al., 2023, Bianchi et al., 2024, Huang et al., 2024, Xia et al., 2024, Abdelnabi et al., 2024], becoming a novel and challenging problem. To our knowledge, there are currently no studies that construct a benchmark to evaluate LLMs’ ability to achieve different objectives of political consensus, but there are some works that have explored LLMs for democratic deliberation and benchmarks in political settings. Below, we will introduce these two aspects separately.
@@
The powerful text generation and information processing capabilities of LLMs have led some studies to explore how they can accelerate the process of democratic deliberation. Konya et al. [2023] design a pipeline allowing LLMs to participate in every stage of democratic elections, aiding in extracting and summarizing complex texts to improve decision-making efficiency. Fish et al. [2023] utilizes LLMs’ generative abilities to synthesize a set of opinions most satisfactory to the majority based on survey results about chatbot personalization. Small et al. [2023] apply LLMs to the deliberation platform Polis [Small et al., 2021], finding that LLMs enhance efficiency but still pose unresolved risks. Bakker et al. [2022], Tessler et al. [2024] fine-tune LLMs to repeatedly generate and refine statements representing a group’s collective stances on social or political issues.
@@
LLMs have been widely applied to political science tasks [Li et al., 2024]. However, political science covers a wide range of research questions, resulting in diverse benchmarks. Kornilova and Eidelman [2019], Arregui and Perarnaud [2022], Klüver et al. [2023], Shu et al. [2024] provide data on texts and the ideologies of their associated political parties, which are used for semantic analysis of texts covering different ideologies. Garzia et al. [2017], Vamvas and Sennrich [2020] extensively collect public comments on various political issues in Europe to study the positioning and classification of political positions. Kornilova and Eidelman [2019], Shu et al. [2024], Arregui and Perarnaud [2022] provide a large collection of legal text data, facilitating research in the generation and summarization of legal documents. POLCA [Moghimifar et al., 2024] collects party statements from several European countries to evaluate whether LLMs can determine if a statement is likely to appear in the final agreement. Stammbach et al. [2024], Chalkidis and Brandl [2024], Batzner et al. [2024] investigate whether LLMs have intrinsic political bias and explore the impact of fine-tuning and prompting on their political stance. Liang et al. [2025] constructs a benchmark based on the United Nations resolution process to evaluate whether LLMs can accurately capture the political stances of member states, simulate voting, and emulate delegate speeches. Although these works offer benchmarks for political science research, their focus is not on studying the ability of LLMs to achieve the objectives of political consensus.
@@ 3.1 Data Collection Procedure
We conduct a large-scale scraping and combine data sourced from the official website of the European Parliament [^1] https://www.europarl.europa.eu , HowTheyVote [^2] https://howtheyvote.eu , and the VoteWatch Europe dataset [HIX et al., 2022], to obtain a comprehensive collection of parliamentary records from the European Parliament spanning a 13-year period from 2009 to 2022. This dataset covers three parliamentary terms of the 7th, 8th, and 9th, as well as includes detailed information on issues, topics, debates, resolutions, and votes.
Unlike previous datasets that were also collected from the European Parliament or political parties [Koehn, 2005, HIX et al., 2022, Chalkidis and Brandl, 2024, Moghimifar et al., 2024], we (1) do not just scrape a single aspect of the parliamentary process, such as debates [Chalkidis and Brandl, 2024] or votes [HIX et al., 2022], but instead collect all information corresponding to each issue from different sources separately, further aligning and integrating them more comprehensively, including information on issues, topics, debates, resolutions, and votes. (2) We perform additional cleaning and post-processing on the data to enhance its quality and readability. (3) The cleaned voting and resolution data can serve as the basis for our open-ended evaluation, allowing further verification of whether our designed evaluation framework aligns with real-world voting outcomes. These contributions not only enhance the quality and diversity of our data but also allow the data to transcend the scope of a single task (such as being used solely for text translation [Koehn, 2005]) and further enable the construction of various complex political tasks and scenarios in PoliCon. We will introduce them one by one as follows:
@@
We first match the URL provided for each issue’s voting information in the VoteWatch Europe dataset with the corresponding issue URLs on the European Parliament’s official website and HowTheyVote. This allows us to obtain the issue and resolution content corresponding to each voting record. We further match the resolution with the debate URL on the European Parliament’s website using the issue name, enabling us to scrape the corresponding debate information. In this way, we obtain 30,698 raw parliamentary records. However, since many records were incomplete or duplicated, we further refine the data, retaining only those where the final vote was confirmed to be finished and all information was complete. The detailed filtering steps are provided in Appendix A.1. Furthermore, referring to the topics defined in the VoteWatch Europe dataset [HIX et al., 2022], we classify all these complete data records into 5 coarse- and 19 fine-grained topics (detailed in Figure 2), such as “culture & education”, “agriculture”, “international trade”, etc. Through this approach, we integrate different pieces of information on the same issue from various sources, ultimately selecting 2,225 complete, high-quality raw data entries, ensuring that each data entry contains a quintuple of raw information: (issue, topic, debates, resolution, votes).
@@
To address raw data redundancy, we employ DeepSeek-R1 [Guo et al., 2025] and rule-based methods for data cleaning. DeepSeek-R1 is used to summarize the relevant background of the issue, select party stances from debates, and remove redundant content from the resolution. Rule-based methods are then applied to randomly perform synonym replacement to diversify the stance data. Voting data is processed by matching each member with their party and calculating party voting results by rounding down the proportion of MEPs within the party who voted in favor to an integer between 0 and 9. This results in cleaned sextuples of (issue, topic, background, stances, resolution, votes) containing relevant party information. Further details of the data cleaning procedure, prompts, and cases for qualification can be found in Appendix A.6, Appendix C.1, and Appendix D.
@@ 3.2 Task Settings
The core requirement of conflict is to have a different number of stakeholders with various positions on each issue. There are clear differences in political positions among the parties in the European Parliament [McElroy and Benoit, 2007, Proksch and Slapin, 2010, McElroy and Benoit, 2012]. To demonstrate this point more obviously, we randomly sample 200 stance data points from each party during the 7th and 8th parliamentary terms. We then use OpenAI’s text-embedding-003-small [OpenAI, 2024] model to map each party’s stances into a semantic representation space, and employ Principal Component Analysis (PCA) [Wold et al., 1987] to visualize this information. As shown in Figure 3, the stances of each party form distinct clusters in the semantic space, with significant differences (detailed in Appendix B.2). We further design three different settings for the number of participating parties in PoliCon: 2, 4, and 6. For each issue, we select the corresponding number of parties with the highest voting variance to enhance the conflict.
@@
We refer to the collective decision-making procedure in different parliaments and the United Nations Security Council (UNSC) to set three common voting mechanisms, which are: (1) Simple Majority: A resolution needs to be voted through by more than 50% of the parliamentary seats. We define the boolean variable $v\in\{0,1\}$ to indicate whether the resolution will be passed under this voting mechanism. In the setting of simple majority, $v=1$ only if $u\geq 5$. (2) Two-thirds Majority: A resolution needs to be voted through by more than two-thirds of the parliamentary seats. In this setting, $v=1$ only if $u\geq 6.67$. (3) Veto Power: To extend PoliCon beyond the context of the European Parliament, we introduce a veto mechanism. In the UNSC, permanent members have veto power [United Nations, ]. In our setting, the tested LLM needs to generate a resolution that can be passed by a simple majority of MEPs in the parliament and not be rejected by the vetoing party (in favor rate under 60%). In this setting, $v=1$ only if $u\geq 5$ and $u_{k}\geq 6$, where $u_{k}$ is the voting score of the vetoing party. In the actual process of constructing the task, we randomly designate which political party has the veto power.
@@
Political goals indicate when the political consensus is achieved. In PoliCon, we define three different political goals as follows: (1) Passing a Resolution: This is the most common political consensus objective, aimed at finding a consensus resolution that can be passed under a specific power structure and voting mechanism detailed above. (2) Rawlsianism: To study the extent to which LLMs can accommodate the interests of minority groups, following the Rawlsian principle [Rawls, 2017], the political goal in this context is to formulate a resolution that maximizes the benefits for the party with the least benefits. In this setting, $u=\text{min}_{i\in n}(u_{i})$. (3) Utilitarianism: Following the Utilitarian principle [Mill, 2016], the political goal is to formulate a resolution that maximizes the sum of benefits for all parties. Under this setting, $u=\sum\limits_{i=1}^{n}u_{i}$.
@@ 4 Experiments
In this section, we use PoliCon to conduct comprehensive experiments to evaluate six current representative LLMs, specifically on their capability to achieve diverse objectives of political consensus. We selected two close-sourced models: GPT-4o [Hurst et al., 2024] and Gemini-2.5-Flash (Gemini-2.5) [DeepMind, 2024], as well as four open-sourced models from different vendors and with varying parameters: Qwen2.5-32B-Instruct (Qwen2.5-32B), Qwen2.5-72B-Instruct (Qwen2.5-72B) [Yang et al., 2024, Team, 2024], Llama-3.3-70B-Instruct (Llama-3.3-70B) [AI@Meta, 2024], and 671-billion-parameter DeepSeek-V3.1 [DeepSeek-AI, 2024]. All LLMs are set up with standardized inference settings, including a temperature of 0.7 and top-p sampling of 0.95. For Gemini-2.5 and DeepSeek-V3.1, we use their thinking versions.

### v3

@@ Abstract
Achieving political consensus is crucial yet challenging for the effective functioning of social governance. However, although frontier AI systems represented by large language models (LLMs) have developed rapidly in recent years, their capabilities in this scope are still understudied. In this paper, we introduce PoliCon, a novel benchmark constructed from 2,225 high-quality deliberation records of the European Parliament over 13 years, ranging from 2009 to 2022, to evaluate the ability of LLMs to draft consensus resolutions based on divergent party positions under varying collective decision-making contexts and political requirements. Specifically, PoliCon incorporates four factors to build each task environment for finding different political consensus: specific political issues, political goals, participating parties, and power structures based on seat distribution. We also developed an evaluation framework based on social choice theory for PoliCon, which simulates the real voting outcomes of different political parties to assess whether LLM-generated resolutions meet the requirements of the predetermined political consensus. Our experimental results demonstrate that even state-of-the-art models remain undersatisfied with complex tasks like passing resolutions by a two-thirds majority and addressing security issues, while uncovering their inherent partisan biases and revealing some behaviors LLMs show to achieve the consensus, such as prioritizing the stance of the dominant party instead of uniting smaller parties, which highlights PoliCon’s promise as an effective platform for studying LLMs’ ability to promote political consensus. The code is released at  PoliCon.
@@ 1 Introduction
One of the fundamental prerequisites for effective social governance is establishing political consensus across diverse stakeholders (Prothro and Grigg, 1960; Lijphart and others, 1999; Huckfeldt et al., 2004; Rawls, 2020). From infrastructure development to welfare policies, consensus-building underpins the legitimacy (Cohen, 2005) and implementation of collective decisions (Citrin, 2001; Potapchuk and Crocker, 2017; Shehu, 2017). Yet, in pluralistic societies, conflicting values, power dynamics, and issue complexity render this process exceptionally challenging (Raiffa, 1982; Ehtamo et al., 1999; Susskind et al., 1999; Baker and Azher, 2024). While large language models (LLMs) have shown promise in facilitating group discussions (Chiang et al., 2024), supporting democratic deliberation (Small et al., 2023; Fish et al., 2023; Tessler et al., 2024; Jarrett et al., 2025), resolving regional conflicts (Konya et al., 2025), and analyzing ideological stances (Chen et al., 2024; Kim et al., 2025), their capacity to find consensus in real and complex political scenarios remains underexplored. This gap raises a critical question: Can LLMs bridge divides among divergent stakeholders and achieve the objectives of different political consensus in real-world settings?
@@
Specifically, PoliCon has designed four adjustable factors to build different task environments, which are: (1) Political issues: the political problems to be discussed and their topic classification, (2) Political goals: the criteria for achieving the corresponding political consensus, (3) Participating parties: different numbers of involving stakeholders with varying stances, and (4) Power structures: differences in influence and discourse power of each party due to their number of seats. By combining these settings, we have constructed a total of 28,620 detailed scenarios. To assess whether LLM-generated resolutions meet the corresponding political goals, we further developed an open-ended evaluation framework in PoliCon based on the social choice theory (Sen, 1986; Kelly, 2013). Through our experiments, we have verified its strong capability to simulate the real voting results, thereby allowing the effective evaluation in PoliCon (subsection 4.1).
@@ 2 Related Work
Achieving political consensus, due to its realistic and complex scenarios, the conflict of stances and values, and the need to consider diverse power structures, differs from existing works that primarily consider conversational grounding (Udagawa and Aizawa, 2019; 2020; 2021; Mitsuda et al., 2022; Mohapatra et al., 2024) and game-theoretic bargaining (Lewis et al., 2017; Paquette et al., 2019; Zhou et al., 2023; Bianchi et al., 2024; Huang et al., 2024; Xia et al., 2024; Abdelnabi et al., 2024), becoming a novel and challenging problem. To our knowledge, there are currently no studies that construct a benchmark to evaluate LLMs’ ability to achieve different objectives of political consensus, but there are some works that have explored LLMs for democratic deliberation and benchmarks in political settings. Below, we will introduce these two aspects separately.
@@
The powerful text generation and information processing capabilities of LLMs have led some studies to explore how they can accelerate the process of democratic deliberation. Konya et al. (2023) design a pipeline allowing LLMs to participate in every stage of democratic elections, aiding in extracting and summarizing complex texts to improve decision-making efficiency. Fish et al. (2023) utilizes LLMs’ generative abilities to synthesize a set of opinions most satisfactory to the majority based on survey results about chatbot personalization. Small et al. (2023) apply LLMs to the deliberation platform Polis (Small et al., 2021), finding that LLMs enhance efficiency but still pose unresolved risks. Bakker et al. (2022); Tessler et al. (2024) fine-tune LLMs to repeatedly generate and refine statements representing a group’s collective stances on social or political issues.
@@
LLMs have been widely applied to political science tasks (Li et al., 2024). However, political science covers a wide range of research questions, resulting in diverse benchmarks. Kornilova and Eidelman (2019); Arregui and Perarnaud (2022); Klüver et al. (2023); Shu et al. (2024) provide data on texts and the ideologies of their associated political parties, which are used for semantic analysis of texts covering different ideologies. Garzia et al. (2017); Vamvas and Sennrich (2020) extensively collect public comments on various political issues in Europe to study the positioning and classification of political positions. Kornilova and Eidelman (2019); Shu et al. (2024); Arregui and Perarnaud (2022) provide a large collection of legal text data, facilitating research in the generation and summarization of legal documents. POLCA (Moghimifar et al., 2024) collects party statements from several European countries to evaluate whether LLMs can determine if a statement is likely to appear in the final agreement. Stammbach et al. (2024); Chalkidis and Brandl (2024); Batzner et al. (2024) investigate whether LLMs have intrinsic political bias and explore the impact of fine-tuning and prompting on their political stance. Liang et al. (2025) constructs a benchmark based on the United Nations resolution process to evaluate whether LLMs can accurately capture the political stances of member states, simulate voting, and emulate delegate speeches. Although these works offer benchmarks for political science research, their focus is not on studying the ability of LLMs to achieve the objectives of political consensus.
@@ 3.1 Data Collection Procedure
We conduct a large-scale scraping and combine data sourced from the official website of the European Parliament [^1] https://www.europarl.europa.eu , HowTheyVote [^2] https://howtheyvote.eu , and the VoteWatch Europe dataset (HIX et al., 2022), to obtain a comprehensive collection of parliamentary records from the European Parliament spanning a 13-year period from 2009 to 2022. This dataset covers three parliamentary terms of the 7th, 8th, and 9th, as well as includes detailed information on issues, topics, debates, resolutions, and votes.
Unlike previous datasets that were also collected from the European Parliament or political parties (Koehn, 2005; HIX et al., 2022; Chalkidis and Brandl, 2024; Moghimifar et al., 2024), we (1) do not just scrape a single aspect of the parliamentary process, such as debates (Chalkidis and Brandl, 2024) or votes (HIX et al., 2022), but instead collect all information corresponding to each issue from different sources separately, further aligning and integrating them more comprehensively, including information on issues, topics, debates, resolutions, and votes. (2) We perform additional cleaning and post-processing on the data to enhance its quality and readability. (3) The cleaned voting and resolution data can serve as the basis for our open-ended evaluation, allowing further verification of whether our designed evaluation framework aligns with real-world voting outcomes. These contributions not only enhance the quality and diversity of our data but also allow the data to transcend the scope of a single task (such as being used solely for text translation (Koehn, 2005)) and further enable the construction of various complex political tasks and scenarios in PoliCon. We will introduce them one by one as follows:
@@
We first match the URL provided for each issue’s voting information in the VoteWatch Europe dataset with the corresponding issue URLs on the European Parliament’s official website and HowTheyVote. This allows us to obtain the issue and resolution content corresponding to each voting record. We further match the resolution with the debate URL on the European Parliament’s website using the issue name, enabling us to scrape the corresponding debate information. In this way, we obtain 30,698 raw parliamentary records. However, since many records were incomplete or duplicated, we further refine the data, retaining only those where the final vote was confirmed to be finished and all information was complete. The detailed filtering steps are provided in Appendix A.1. Furthermore, referring to the topics defined in the VoteWatch Europe dataset (HIX et al., 2022), we classify all these complete data records into 5 coarse- and 19 fine-grained topics (detailed in Figure 2), such as “culture & education”, “agriculture”, “international trade”, etc. Through this approach, we integrate different pieces of information on the same issue from various sources, ultimately selecting 2,225 complete, high-quality raw data entries, ensuring that each data entry contains a quintuple of raw information: (issue, topic, debates, resolution, votes).
@@
To address raw data redundancy, we employ DeepSeek-R1 (Guo et al., 2025) and rule-based methods for data cleaning. DeepSeek-R1 is used to summarize the relevant background of the issue, select party stances from debates, and remove redundant content from the resolution. Rule-based methods are then applied to randomly perform synonym replacement to diversify the stance data. Voting data is processed by matching each member with their party and calculating party voting results by rounding down the proportion of MEPs within the party who voted in favor to an integer between 0 and 9. This results in cleaned sextuples of (issue, topic, background, stances, resolution, votes) containing relevant party information. Further details of the data cleaning procedure, prompts, and cleaning qualification can be found in Appendix A.6, Appendix C.1, and Appendix D.
@@ 3.2 Task Settings
The core requirement of conflict is to have a different number of stakeholders with various positions on each issue. There are clear differences in political positions among the parties in the European Parliament (McElroy and Benoit, 2007; Proksch and Slapin, 2010; McElroy and Benoit, 2012). To demonstrate this point more obviously, we randomly sample 200 stance data points from each party during the 7th and 8th parliamentary terms. We then use OpenAI’s text-embedding-003-small (OpenAI, 2024) model to map each party’s stances into a semantic representation space, and employ Principal Component Analysis (PCA) (Wold et al., 1987) to visualize this information. As shown in Figure 3, the stances of each party form distinct clusters in the semantic space, with significant differences (detailed in Appendix B.2). We further design three different settings for the number of participating parties in PoliCon: 2, 4, and 6. For each issue, we select the corresponding number of parties with the highest voting variance to enhance the conflict.
@@
We refer to the collective decision-making procedure in different parliaments and the United Nations Security Council (UNSC) to set three common voting mechanisms, which are: (1) Simple Majority: A resolution needs to be voted through by more than 50% of the parliamentary seats. We define the boolean variable $v\in\{0,1\}$ to indicate whether the resolution will be passed under this voting mechanism. In the setting of simple majority, $v=1$ only if $u\geq 5$. (2) Two-thirds Majority: A resolution needs to be voted through by more than two-thirds of the parliamentary seats. In this setting, $v=1$ only if $u\geq 6.67$. (3) Veto Power: To extend PoliCon beyond the context of the European Parliament, we introduce a veto mechanism. In the UNSC, permanent members have veto power (United Nations, ). In our setting, the tested LLM needs to generate a resolution that can be passed by a simple majority of MEPs in the parliament and not be rejected by the vetoing party (in favor rate over 60%). In this setting, $v=1$ only if $u\geq 5$ and $u_{k}\geq 6$, where $u_{k}$ is the voting score of the vetoing party. In the actual process of constructing the task, we randomly designate which political party has the veto power.
@@
Political goals indicate when the political consensus is achieved. In PoliCon, we define three different political goals as follows: (1) Passing a Resolution: This is the most common political consensus objective, aimed at finding a consensus resolution that can be passed under a specific power structure and voting mechanism detailed above. (2) Rawlsianism: To study the extent to which LLMs can accommodate the interests of minority groups, following the Rawlsian principle (Rawls, 2017), the political goal in this context is to formulate a resolution that maximizes the benefits for the party with the least benefits. In this setting, $u=\text{min}_{i\in n}(u_{i})$. (3) Utilitarianism: Following the Utilitarian principle (Mill, 2016), the political goal is to formulate a resolution that maximizes the sum of benefits for all parties. Under this setting, $u=\sum\limits_{i=1}^{n}u_{i}$.
@@ 4 Experiments
In this section, we use PoliCon to conduct comprehensive experiments to evaluate six current representative LLMs, specifically on their capability to achieve diverse objectives of political consensus. We selected two close-sourced models: GPT-4o (Hurst et al., 2024) and Gemini-2.5-Flash (Gemini-2.5) (DeepMind, 2024), as well as four open-sourced models from different vendors and with varying parameters: Qwen2.5-32B-Instruct (Qwen2.5-32B), Qwen2.5-72B-Instruct (Qwen2.5-72B) (Yang et al., 2024; Team, 2024), Llama-3.3-70B-Instruct (Llama-3.3-70B) (AI@Meta, 2024), and 671-billion-parameter DeepSeek-V3.1 (DeepSeek-AI, 2024). All LLMs are set up with standardized inference settings, including a temperature of 0.7 and top-p sampling of 0.95. For Gemini-2.5 and DeepSeek-V3.1, we use their thinking versions.
To clearly demonstrate the model’s capabilities, we also introduce two rule-based baselines: Random and Greedy. The Random method randomly selects the stance of a party as the final resolution, while the Greedy method selects the stance of the party with the largest number of seats as the resolution.